OpenStudy (anonymous):

I am trying to graph exponential functions and I am a mom that is confused. The equation is y=-2 x 5^x. Can anyone help

OpenStudy (whpalmer4):

Yes, I can help you. What is your first question? Do you have any idea about how to proceed?

OpenStudy (anonymous):

You need to make a graph for the points on the graph

OpenStudy (whpalmer4):

Okay, what are you going to use for points? The equation is \[y = -2*5^x\]correct?

OpenStudy (anonymous):

Starting at -1

OpenStudy (whpalmer4):

Let's say we want to go through x = 3. I would make a table, with the left column being the values of x (-1, 0, 1, 2, 3) and the right column being the corresponding values of y, which you find by plugging x into the equation. For example, next to x = 3, you would have \(y = -2*5^3 = -2*125 = -250\). When your table is complete, then you can plot those points and draw a smooth curve through them.

OpenStudy (anonymous):

I was working along with and got the same thing for 3. I appreciate the help.

OpenStudy (whpalmer4):

Here's what my graph ended up looking like.

OpenStudy (anonymous):

Yep thats what we got